OpenStudy (anonymous):

x+9x-6x+4x=20 help??

OpenStudy (anonymous):

its simple addition. dont look at the x's, look at the number infront of the x's. and if the an x is alone by itself, from now on its considered a 1 infront of the x. so first add all the numbers infront of the x's. once you have dont that type a respond.

OpenStudy (anonymous):

x+9x-6x+4x=20 (x+9x)-6x+4x=20 10x-6x+4x =20 4x+4x =20 8x = 20 x = 20/8 x = 5/2

OpenStudy (anonymous):

okay looks like you are AFK. so once you have done adding/subtracting the numbers. 1+9-6+4=8. so you get 8, but not you have to account for that x. so put the x after the 8. so 8x=20. and now to get x alone we divide by 8 from both sides  20 over 8 can be simplified to 5/2. so therefore x is equal to 5 over 2